OpenStudy (itiaax):

*MEDAL WILL BE AWARDED* The average of 5 numbers is 12. If a 6th number is added, what is the average?

OpenStudy (zehanz):

You'll have to know what number is added!

OpenStudy (itiaax):

Really? :(

OpenStudy (anonymous):

i was going to say, without the 6th number you cant do it

OpenStudy (zehanz):

Try it with an example: The average of 12, 12, 12, 12 and 12 (5 numbers) is ...12! Now add 12. Average is still 12. Instead, add 60. Average is 120/6=20. Makes a lot of difference.

OpenStudy (anonymous):

If "x" is your 6th number, your average is:\[\frac{ 60 + x }{ 6 }\]This gives a real and useful answer, but of necessity has to be left in the form of a formula.

OpenStudy (anonymous):

In case that Latex didn't come through: (x + 60) / 6

OpenStudy (anonymous):

"60" is from: (x1 + x2 + x3 + x4 + x5)/5 = 12, so the sum is 60 for the first 5 numbers.

OpenStudy (anonymous):

Making sense now, @iTiaax ?

OpenStudy (itiaax):

Yes :) Thank you so much

OpenStudy (anonymous):

You're quite welcome!

OpenStudy (anonymous):

And thx for the recognition! @iTiaax ! Good luck in all of your studies!

OpenStudy (itiaax):

No problem :) Good luck to you too!